OpenStudy (anonymous):

someone PLEASE help me. It says recipe A makes 5 dinner rolls using 1 cup of flour. Recipe B makes 24 rolls using 7 1/2 c of flour. Recipe C makes 45 rolls using 10 c of flour. which recipe requires the most flour per roll?

OpenStudy (studygurl14):

Recipe A: 5 dinner rolls with 1 cup of flour 1 cup / 5 rolls equals 1/5 cup per roll Recipe B: 24 rolls using 7.5 cups of flour 7.5 / 24 = 5/16 cup per roll Recipe C: 45 rolls, 10 cups of flour 10/45 = 2/9 cups per roll 1/5 = 0.2 5/16 = 0.3125 2/9 = 0.2222222... So...which recipe requires the most flour per roll?

OpenStudy (anonymous):

recipe B?

OpenStudy (anonymous):

but how did you get 5/16?

OpenStudy (studygurl14):

7.5 / 24 = 5/16

OpenStudy (studygurl14):

See this part: "Recipe B: 24 rolls using 7.5 cups of flour 7.5 / 24 = 5/16 cup per roll "

OpenStudy (anonymous):

yeah but 7.5 divided by 24 is 0.3125 and 7.5 times 24 is 180 so?

OpenStudy (studygurl14):

7.5/24 = 5/16 = 0.3125

OpenStudy (anonymous):

is that simplified or what?

OpenStudy (anonymous):

like how did you get 5/16?

OpenStudy (studygurl14):

oh...yeah i don't like decimals infractions so i "edited" it